OpenStudy (anonymous):

can anybody help me on - how to find the fastest way to check if a large number is perfect square or not?

OpenStudy (tkhunny):

0^2 = 0 1^2 = 1 2^2 = 4 3^2 = 9 4^2 = _6 5^2 = _5 6^2 = _6 7^2 = _9 8^2 = _4 9^2 = _1 It appears, if it ends in 0, 1, 4, 5, 6, or 9, it COULD BE a perfect square. Ends with 2, 3, 7, or 8? Don't consider looking any more. It's not a perfect square. What other ways can you find that might help determine a particular value?

OpenStudy (anonymous):

yes, I got this. actually, it a huge number of 20 digits. so, I needed a way to find if this is the perfect square or not. your answer solved my problem. Thanks buddy !

OpenStudy (anonymous):

acually, my question was:

OpenStudy (anonymous):

All the nine digits are arranged here so as to form four square numbers : 9, 81, 324, 576 how would you put them together so as to form a single smallest possible square number and a single largest possible square number?

OpenStudy (dan815):

hmm

OpenStudy (dan815):

ok not that!

OpenStudy (dan815):

but i am thinking eulers method could be 1 way

OpenStudy (dan815):

i think just eulers formula plugged into a computer will be the fastest way to see if its a sqrt or not, as soon as it settles on a value

OpenStudy (dan815):

you just have to show its converging to a number in the end